# Fine-Tuning Code Language Models to Detect Cross-Language Bugs

### v1

@@ 3.2. Cross-Language Code Identification Tool
Currently, there have been no tools for identifying cross-language code. To fill this gap, we developed CLCFinder specifically designed for the interaction characteristics of Python-C/C++, Java-C/C++, and Python-Java language combinations. We used CLCFinder to collect cross-language code data. As shown in Table 1, it can identify nine cross-language interaction mechanisms across Python, C/C++, and Java, including six between Python and C, one between Python and Java, and two between Java and C/C++.
@@ 3.4. Selection of CodeLMs
Our research goal is to fine-tune CodeLMs on our CLB dataset to detect CLBs. The following are the criteria we employed to select the CodeLMs: Firstly, since we need to fine-tune the CodeLMs, they must be open-source. Open-source CodeLMs ensure the transparency and reproducibility of the research, as well as provide more opportunities for customization and optimization. Secondly, the CodeLMs need to support binary classification. We treat the code bug prediction task as a binary classification problem, i.e., determining whether a code snippet contains a bug. Thirdly, large pre-trained models typically require significant computational resources (such as GPUs or TPUs). Therefore, the CodeLMs that we choose must balance good performance and reasonable resource consumption, not beyond our computational capacity. Lastly, the model should be one of the popular or latest CodeLMs, ensuring that the CodeLMs we use are competitive in code understanding and bug detection tasks. Based on these criteria, we selected 13 CodeLMs for our experiments. Table 2 shows the detailed information about these models. These models have demonstrated outstanding performance in tasks such as code understanding, code generation, code search, and code completion (Feng et al., 2020; Guo et al., 2020; Bai et al., 2023; Hui et al., 2024; Guo et al., 2024c).
@@
Deepseek-Coder-1.3B-base (Guo et al., 2024c) 1.3B 16384 Decoder-only
@@ 3.6. Fine-tuning Environments and Methods for CodeLMs
NVIDIA GeForce
RTX 4090 24G
CodeBERT-base, Graph CodeBERT-base, UniXcoder, Full fine-tuning
@@ 4.2.1. Performance Metrics of CodeLMs
We fine-tuned 13 well-known open-source CodeLMs and evaluated their performance in detecting CLBs, and the detailed results are presented in Table 4. By comparing these results, we found that most CodeLMs show some improvement in the CLB detection task. The detailed comparisons of the model performance metrics before and after fine-tuning are presented in the following.
@@
Table 4. Performance of CodeLMs in detecting CLBs before and after fine-tuning on the CLB dataset without code comments
CodeLM Performance Before Fine-Tuning Performance After Fine-Tuning Performance Change (%)
Acc Prec Rec F1 AUC Acc Prec Rec F1 AUC Acc Prec Rec F1 AUC
CodeBERT-base 0.5009 0.5005 0.8381 0.6268 0.5172 0.6331 0.6022 0.7842 0.6813 0.7131 26.39 20.32 -6.43 8.69 37.88
GraphCodeBERT-base 0.5153 0.5167 0.4730 0.4939 0.5159 0.6517 0.6128 0.8241 0.7029 0.7447 26.47 18.60 74.23 42.32 44.35
UniXcoder-base 0.4946 0.4946 0.4964 0.4955 0.4956 0.7172 0.6839 0.8079 0.7407 0.8088 45.01 38.27 62.75 49.49 63.20
CodeT5-base 0.5018 0.5016 0.5791 0.5376 0.4886 0.6454 0.6350 0.6840 0.6586 0.7238 28.62 26.59 18.11 22.51 48.14
CodeT5p-220M 0.4964 0.4975 0.7122 0.5858 0.4914 0.6364 0.6148 0.7307 0.6678 0.7313 28.20 23.58 2.60 14.00 48.82
CodeT5p-770M 0.4931 0.4674 0.4323 0.4492 0.4915 0.5302 0.5238 0.8779 0.6561 0.5432 7.52 12.07 103.08 46.06 10.52
NatGen 0.5009 0.5007 0.6906 0.5805 0.5088 0.6364 0.6095 0.7594 0.6763 0.7278 27.05 21.73 9.96 16.50 43.04
Deepseek-Coder-1.3B-base 0.5184 0.5051 0.1821 0.2677 0.4704 0.5498 0.5255 0.7063 0.6026 0.6125 6.06 4.04 287.86 125.10 30.21
Granite-3B-Code-base-2k 0.4851 0.4806 0.6135 0.5390 0.4811 0.5516 0.5275 0.6951 0.5998 0.5865 13.71 9.76 13.30 11.28 21.91
StarCoder2-3B 0.5076 0.4987 0.7125 0.5867 0.5156 0.5561 0.5358 0.7124 0.6116 0.5889 9.55 7.44 -0.01 4.24 14.24
CodeQWen1.5-7B 0.5058 0.4957 0.4322 0.4618 0.5062 0.5229 0.5043 0.7580 0.6057 0.5706 3.38 1.73 75.37 31.15 12.72
Qwen2.5_Coder-7B 0.4761 0.4735 0.6062 0.5317 0.4719 0.5238 0.5066 0.5669 0.5350 0.5676 10.02 6.99 -6.48 0.62 20.28
CodeLlama-7b-hf 0.5049 0.4860 0.4200 0.4506 0.5039 0.4979 0.5079 0.5269 0.5172 0.5442 -1.39 4.51 25.45 14.78 8.00
Accuracy: As shown in Table 4, the accuracy of these 13 models improved almost uniformly after fine-tuning. The greatest improvement was seen in UniXcoder-base, with an increase of 45.01%, followed by CodeT5-base with a 28.62% improvement. However, the accuracy of CodeLlama-7b-hf was decreased by 1.39% after fine-tuning, which is the only model with a declined accuracy after fine-tuning. This indicates that fine-tuning effectively enhances the overall prediction accuracy of the models to some extent.
Precision: After fine-tuning, the precision of all models were improved to different degrees. Specifically, the precision of most models significantly improved, particularly UniXcoder-base (with an increase of 38.27%) and Qwen2.5_Coder-7B (with an increase of 34.59%). These improvements suggest that fine-tuning enhanced the model’s ability to identify positive instances, thereby increasing prediction accuracy. However, although the precision of CodeQWen1.5-7B and Deepseek-Coder-1.3B-base also improved, the changes were relatively small, with increases of only 1.73% and 4.04%, respectively.
Recall: Recall reflects the model’s ability to capture positive instances. The recall of these 13 models did not always improve after fine-tuning, and the changes varied significantly. For example, GraphCodeBERT-base improved from 0.1821 before fine-tuning to 0.7063 after fine-tuning, indicating that fine-tuning effectively enhanced the model’s sensitivity in recognizing positive instances. However, some models did not show improvement after fine-tuning, such as Qwen2.5_Coder-7B, which decreased by 6.48% (from 0.6062 before fine-tuning to 0.5669 after fine-tuning).
F1 score: The F1 scores of all CodeLMs were improved after fine-tuning, particularly Deepseek-Coder-1.3B-base (increasing from 0.2677 to 0.6026) and UniXcoder-base (increasing from 0.4955 to 0.7407). This change indicates that fine-tuning not only enhanced the CodeLM’s accuracy and recall but also further optimized its ability to handle the balance between positive and negative classes. Therefore, fine-tuning significantly improved the CodeLM’s overall performance in CLB detection tasks.
AUC: The AUC of all CodeLMs were enhanced after fine-tuning. Specifically, most CodeLMs showed a significant increase in AUC after fine-tuning. For example, UniXcoder-base had the largest improvement in AUC (63.20%), increasing from 0.5956 to 0.8088, while the model with the smallest improvement, CodeLlama-7b-hf, still increased by 8.00% (from 0.5039 to 0.5442). The increase in AUC indicates that fine-tuning improved the CodeLM’s ability to recognize CLB code, particularly in challenging scenarios such as semantically similar code snippets, where fine-tuning helped the model better distinguish between positive and negative samples.

### v2

@@ 3.2. Cross-Language Code Identification Tool
Currently, there have been no tools for identifying cross-language code. To fill this gap, we developed CLCFinder specifically designed for the interaction characteristics of Python-C/C++, Java-C/C++, and Python-Java language combinations. CLCFinder is included in the replication package of this work to support reproducibility. We used CLCFinder to collect cross-language code data. As shown in Table 1, it can identify nine cross-language interaction mechanisms across Python, C/C++, and Java, including six between Python and C, one between Python and Java, and two between Java and C/C++.
The selection of these nine cross-language interaction mechanisms follows the principles of representativeness and practical feasibility. Python, C/C++, and Java have long been among the most widely used PLs according to GitHub statistics (GitHub Octoverse (GitHub Staff, 2025)) and language popularity rankings such as TIOBE (TIOBE Software, 2026), and they frequently coexist in real-world open-source projects. Motivated by this observation, we surveyed and analyzed the common pairwise cross-language interaction mechanisms among these three languages. For each interaction mechanism, we conducted an in-depth analysis of its characteristics and designed corresponding code analysis logic, which was then implemented in the CLCFinder tool. The final set of nine mechanisms was determined based on extensive investigation of documentation, open-source codebases, and real-world projects, revealing that these mechanisms are widely used in practice and exhibit relatively stable implementation patterns, making them amenable to automated identification. It is worth noting that some cross-language interaction mechanisms allow flexible or diverse implementations; therefore, CLCFinder may not capture all possible variants. However, our design prioritizes the correctness of identified results, ensuring that the collected samples indeed involve genuine cross-language code interactions.
@@ 3.4. Selection of CodeLMs
Our research goal is to fine-tune CodeLMs on our CLB dataset to detect CLBs. The following are the criteria we employed to select the CodeLMs: Firstly, since we need to fine-tune the CodeLMs, they must be open-source. Open-source CodeLMs ensure the transparency and reproducibility of the research, as well as provide more opportunities for customization and optimization. Secondly, the CodeLMs need to support binary classification. We treat the code bug prediction task as a binary classification problem, i.e., determining whether a code snippet contains a bug. Thirdly, large pre-trained models typically require significant computational resources (such as GPUs or TPUs). Therefore, the CodeLMs that we choose must balance good performance and reasonable resource consumption, not beyond our computational capacity. Lastly, the model should be one of the popular or latest CodeLMs, ensuring that the CodeLMs we use are competitive in code understanding and bug detection tasks. Based on these criteria, we selected 13 CodeLMs for our experiments. Table 2 shows the detailed information about these models. These models have demonstrated outstanding performance in tasks such as code understanding, code generation, code search, and code completion (Feng et al., 2020; Guo et al., 2020; Bai et al., 2023; Hui et al., 2024; Guo et al., 2024a).
@@
Deepseek-Coder-1.3B-base (Guo et al., 2024a) 1.3B 16384 Decoder-only
@@ 3.6. Fine-tuning Environments and Methods for CodeLMs
NVIDIA GeForce RTX 4090 24G CodeBERT-base, Graph CodeBERT-base, UniXcoder, Full fine-tuning
@@
3.7. Experimental Design for Each Research Question
Before formally introducing the experimental procedure, we first clarify that in all experimental settings of this study, the input context length of the CodeLMs is uniformly limited to no more than 512 tokens, except for the token sequence related experiments in RQ3, where different maximum token lengths are explicitly evaluated. The rationale for this setting can be summarized as follows. (1) This maximum token length is mainly constrained by the hardware resources used in the experiments, especially the available GPU memory. For models based on the Transformer architecture, the memory and computation cost of the self attention mechanism increases approximately with the square of the input sequence length [73]. (2) From the perspective of data distribution, this length limitation is reasonable. As shown in Figure 4, statistics based on the tokenizers of some CodeLMs indicate that most samples (approximately 75 percent) contain no more than 512 tokens after tokenization. Therefore, this limitation does not significantly affect the information completeness of most samples. (3) This setting is also aligned with the architectural constraints of the evaluated models, as 6 out of the 13 CodeLMs used in this study support a maximum context length of 512 tokens (see Table 2).
In the implementation process, the original input is first tokenized. If the generated token sequence length exceeds the predefined upper limit of 512, only the first 512 tokens are kept and the remaining tokens are discarded. If the sequence length is less than or equal to 512, the entire sequence is preserved and fed into the model. This fixed length truncation strategy effectively controls the input size of the model and avoids GPU memory overflow caused by excessively long sequences.
3.7.1. RQ1: Experimental Setup for Baseline and Fine-Tuning on CLB Dataset
In the RQ1 experiment, we first randomly split the CLB dataset into training, validation, and test sets with a ratio of 8:1:1. We then loaded 13 pre-trained CodeLMs using the AutoModelForSequenceClassification interface and adapted all of them to a binary classification task so that the models can produce classification results.
Table 4. Fine-tuning Hyperparameter Settings for Different CodeLMs
Model Acc. Steps Batch Epochs LR WD Warmup
CodeBERT-base 4 32 20 $\times 10^{-4} 1×1$$\times 10^{-2}$ 0.10
GraphCodeBERT-base 4 32 20 $\times 10^{-4} 1×1$$\times 10^{-2}$ 0.10
UniXcoder-base 4 32 20 $\times 10^{-4} 1×1$$\times 10^{-2}$ 0.10
CodeT5-base 16 8 20 $\times 10^{-4} 1×1$$\times 10^{-2}$ 0.10
CodeT5p-220M 16 8 20 $\times 10^{-4} 1×1$$\times 10^{-2}$ 0.10
CodeT5p-770M 32 4 10 $\times 10^{-4} 1×1$$\times 10^{-2}$ 0.05
NatGen 16 8 20 $\times 10^{-4} 1×1$$\times 10^{-2}$ 0.10
DeepSeek-Coder-1.3B-base 16 8 10 $5\times 10^{-5}$ $1\times 10^{-2}$ 0.06
Granite-3B-Code-base-2k 16 4 15 $3\times 10^{-4}$ $1\times 10^{-2}$ 0.06
StarCoder2-3B 16 4 10 $3\times 10^{-4}$ $1\times 10^{-2}$ 0.05
CodeQwen1.5-7B 64 1 10 $3\times 10^{-4}$ $1\times 10^{-2}$ 0.06
Qwen2.5_Coder-7B 64 1 10 $3\times 10^{-4}$ $1\times 10^{-2}$ 0.06
CodeLLama-7b-hf 4 24 10 $\times 10^{-4} 1×1$$\times 10^{-2}$ 0.06
Note: Acc. Steps = Gradient Accumulation Steps; Batch = Per-device Batch Size; Epochs = Epochs Size; LR = Learning Rate; WD = Weight Decay; Warmup = Warmup Ratio.
In addition, we introduced a general LLM as an inference-based reference baseline for comparison. Considering the tradeoff between performance and inference cost, we selected GPT-4o-mini as the baseline model. This model is only used for inference on the test set and is not involved in any training or fine-tuning process. The prompt used for inference is shown in Table 5.
Next, we fine-tuned the 13 CodeLMs on the CLB dataset. The training set is used for model training, and the validation set is used for model selection and hyperparameter tuning. After fine-tuning, we evaluated the fine-tuned models on the test set to assess their CLB detection performance. The hyperparameters used for fine-tuning are shown in Table 4. These hyperparameters are chosen after many experiments and adjustments to ensure stable and representative performance for all CodeLMs. All CodeLMs are trained using the AdamW (Loshchilov and Hutter, 2017) optimizer with a linear learning rate scheduler and a warmup strategy. The specific learning rate schedule and warmup settings vary across CodeLMs, and the details are reported in Table 4. It is important to note that the hyperparameters used for each model are kept consistent across all experiments in this study. Specifically, for RQ1, RQ2, RQ3, and RQ4, the same model always adopts an identical set of hyperparameters. This design avoids potential interference caused by hyperparameter variations and ensures fair and comparable results across different experiments.
Table 5. GPT-4o-mini Prompt Used for CLB Detection
Role Prompt
System You are an expert software bug detection system specialized in cross-language bugs. A cross-language bug refers to a defect that occurs at the interaction boundary between two different programming languages (e.g., JNI, ctypes, FFI, native extensions), caused by incorrect assumptions or mismatches across languages. These bugs typically involve, but are not limited to: Incorrect data type mapping or size mismatch Memory management issues across language boundaries Improper pointer or reference handling Incorrect function signatures or calling conventions Resource lifecycle mismatches between languages Given the following code snippet, determine whether it contains a cross-language bug. Focus only on bugs that originate from or are triggered by cross-language interactions. Do NOT consider purely single-language bugs. Respond with exactly one word: YES if the code contains a cross-language bug NO if it does not Do not provide explanations or any additional text.
User Please analyze this function code for cross-language bugs: {code}
3.7.2. RQ2: Experimental Setup for Transfer from Single-Language Bug Datasets
Tables 6 and 7 provide the specific details of Java and Python code in the CodeNet and CVEfixes datasets, respectively. The following are the methods we used to process these two datasets. Since our CLB dataset is balanced (with equal numbers of bug and clean code samples), we processed the CVEfixes and CodeNet datasets to ensure fair experimental results. For the CVEfixes dataset, we used all bug samples. As there are more clean samples than bug samples, we randomly selected an equal number of clean samples to match the bug count, resulting in 3,102 pairs of Java bug and clean samples, and 2,775 pairs of Python bug and clean samples. Given the large size of the CodeNet dataset, which contains over 3 million Java and Python code samples, utilizing the entire dataset for fine-tuning would be computationally expensive and impractical. To ensure consistency with our CLB dataset, which contains 11,126 samples, we constructed a subset of the CodeNet dataset for our fine-tuning experiments. Specifically, we first balanced the dataset by randomly selecting an equal number of clean samples to match the number of bug samples. Subsequently, we combined Java and Python samples by class (i.e., merging Java bug samples with Python bug samples, and Java clean samples with Python clean samples), and then randomly selected 10,000 samples each from the bug and clean data to create the final dataset for fine-tuning.
Table 6. Details of the CodeNet
Language Category Count* Sum
Java Buggy 305,759 624,479
Clean 318,720
Python Buggy 1,387,527 2,959,324
Clean 1,571,797
* Only count the results after deduplication.
Table 7. Details of the CVEfixes (the part of Python and Java)
Language Category Count Sum
Java Buggy 3,102 8,387
Clean 5,285
Python Buggy 2,775 7,568
Clean 4,793
Specifically, we first randomly split each dataset into training, validation, and test sets with a ratio of 8:1:1. We then selected five CodeLMs that achieved the best performance on the CLB detection task in RQ1 as the subjects of this study. Next, we fine-tuned these five CodeLMs on the training sets of CVEfixes and CodeNet, respectively, using the corresponding validation sets for model selection and hyperparameter tuning. After fine-tuning, the resulting CodeLMs were directly evaluated on the CLB test set to analyze the extent to which single-language bug knowledge can be transferred to CLB detection. In addition, we used the results obtained by directly fine-tuning the same five CodeLMs on the CLB dataset and evaluating them on the CLB test set as a baseline. This baseline setting is consistent with the experimental setup in RQ1 and serves as a reference for comparing the impact of different fine-tuning datasets on CLB detection performance.
3.7.3. RQ3: Experimental Setup for Data Scale and Token Length Analysis
In RQ3, we analyzed the impact of data scale and input sequence length on model performance in the CLB detection task. Consistent with RQ2, we selected the same five CodeLMs that achieved the best performance in RQ1 as the models under study.
Data scale analysis. Based on the full CLB dataset, we constructed five subsets that contain 20 percent, 40 percent, 60 percent, 80 percent, and 100 percent of the original dataset. Each subset is randomly sampled from the full CLB dataset, while maintaining a similar dataset structure to the original dataset, such as language combinations and label distribution, with a balanced 1:1 ratio between positive and negative samples in each subset. For each subset, the data are randomly split into training, validation, and test sets with a ratio of 8:1:1. For example, when using 60 percent of the full CLB dataset, the subset contains 6,676 code samples, with 5,340 samples for training, 668 samples for validation, and 668 samples for testing. When using the full dataset, all 11,126 code samples are included, with 8,902 samples for training, 1,112 samples for validation, and 1,112 samples for testing. We fine-tuned the models on the corresponding training sets and evaluated their performance on the test sets to analyze how model performance changes under different data scale settings.
Token sequence length analysis. For the input length analysis, we used the full CLB dataset and keep the same splits for the training, validation, and test sets. During model training, validation, and testing, we limited the maximum input token length to 128, 256, 384, and 512. When the input sequence length exceeds the predefined limit, a consistent truncation strategy is applied, where only tokens within the maximum length are kept and the remaining tokens are discarded. Through this experimental setup, we analyze the impact of different input length constraints on CLB detection performance.
3.7.4. RQ4: Experimental Setup for Evaluating the Effect of Code Comments
The experimental setup for RQ4 is consistent with that of RQ1, with the only difference being the dataset preprocessing strategy. In this RQ, we used the original code dataset without removing code comments for CodeLMs fine-tuning and testing, in order to evaluate the impact of code comment information on CLB detection performance.
@@ 4.1. Overview of the Obtained Dataset
While the above analyses characterize the CLB dataset from a structural and statistical perspective, they do not fully reflect the semantic nature of CLBs themselves. To provide a more comprehensive overview of the obtained dataset, we further analyze the observable symptoms and underlying root causes of CLBs.
In addition, we conduct a systematic analysis of the symptoms and root causes of CLBs in the CLB dataset, following the methodology proposed by Yang et al (Yang and Cai, 2025). Since manually inspecting all bug instances is time-consuming, we randomly sample 400 bug cases from the entire dataset for analysis. This sample size is statistically significant with a 95% confidence level and a 5% margin of error. All sampled bug cases are manually examined and categorized by the first and second authors, and any disagreements are resolved through discussion until a consensus is reached. During symptom and root cause classification, we adopt the taxonomy from Yang et al. and make minor adjustments to several categories to better fit the characteristics of the CLB dataset. The detailed distributions of symptoms and root causes are illustrated in Figures 5 and 6, respectively.
Overall, CLBs exhibit relatively concentrated distributions in both their root causes and observable symptoms, while also demonstrating clear characteristics of cross-language interactions. These distributions offer an intuitive view of how CLBs are triggered and how they manifest at runtime.
Figure 5. Distribution of Cross-Language Bug Symptoms in the CLB Dataset
Regarding observable symptoms, CLBs tend to manifest as non-fatal yet hard-to-detect runtime anomalies. Behavioral deviation is the most frequently observed symptom, indicating that many CLBs do not lead to crashes or explicit error reports, but instead appear as unexpected changes in program behavior or output. Error or warning messages and incorrect results or outputs are also commonly observed, showing that some CLBs are exposed through explicit but non-terminating signals. In contrast, overt failures such as crashes and memory leaks occur in relatively fewer cases, suggesting that CLBs are often more hidden and therefore more challenging to debug and localize. Additionally, it should be noted that the “Other” category accounts for approximately 10% of the sampled instances. This category includes various uncommon bug symptoms that cannot be classified into the existing major categories, such as (but not limited to) logging and observability issues, API deprecation mismatches at language boundaries, and metric reporting failures. These types are grouped into “Other” mainly due to their very small sample sizes. Furthermore, this category also includes cases where the specific symptom cannot be determined due to insufficient descriptions in the original commit messages and issue reports.
Figure 6. Distribution of Root Causes of CLBs in the CLB Dataset
From the perspective of root causes, issues directly related to cross-language integration dominate the dataset. Among them, cross-language integration errors constitute the largest category, indicating that many bugs originate from inconsistencies at language boundaries, such as cross-language interface invocations and misalignment of data and control flows. In addition, errors introduced by system evolution and refactoring also account for a substantial number of cases, suggesting that implicit cross-language assumptions are easily violated during the long-term evolution of multilingual systems. Meanwhile, data type and cross-language representation errors, as well as performance and optimization related errors, are also observed in non-negligible quantities, reflecting the complexity of data representation and execution semantics in cross-language systems.
The statistical analysis of CLB symptoms and root causes indicate that bugs in the CLB dataset exhibit strong cross-language characteristics. Their root causes are mainly concentrated at language boundaries, system evolution, and data representation, while their symptoms are more likely to appear as behavioral deviations rather than immediate failures. This analysis further highlights the inherent complexity of cross-language bugs in real-world software systems and provides empirical support for the necessity of dedicated modeling and research on CLBs.
@@ 4.2.1. Performance Metrics of CodeLMs
We fine-tuned 13 well-known open-source CodeLMs and evaluated their performance in detecting CLBs. In addition, we include GPT-4o-mini as an inference-only baseline, which is evaluated on the same test set using a prompt-based reasoning strategy, without any task-specific fine-tuning. The detailed results are presented in Table 8, note that the entries for GPT-4o-mini only report inference performance. The detailed comparisons of the fine-tuned CodeLMs performance metrics, alongside the GPT-4o-mini baseline, are presented in the following.
@@
Table 8. Performance of CodeLMs in detecting CLBs after fine-tuning on the CLB dataset without code comments
CodeLM Acc Prec Rec F1 AUC
GPT-4o-mini (Baseline) 0.5196 0.5190 0.5357 0.5272 0.5196
CodeBERT-base 0.6331 0.6022 0.7842 0.6813 0.7131
GraphCodeBERT-base 0.6517 0.6128 0.8241 0.7029 0.7447
UniXcoder-base 0.7172 0.6839 0.8079 0.7407 0.8088
CodeT5-base 0.6454 0.6350 0.6840 0.6586 0.7238
CodeT5p-220M 0.6364 0.6148 0.7307 0.6678 0.7313
CodeT5p-770M 0.5302 0.5238 0.8779 0.6561 0.5432
NatGen 0.6364 0.6095 0.7594 0.6763 0.7278
Deepseek-Coder-1.3B-base 0.5498 0.5255 0.7063 0.6026 0.6125
Granite-3B-Code-base-2k 0.5516 0.5275 0.6951 0.5998 0.5865
StarCoder2-3B 0.5561 0.5358 0.7124 0.6116 0.5889
CodeQWen1.5-7B 0.5229 0.5043 0.7580 0.6057 0.5706
Qwen2.5_Coder-7B 0.5238 0.5066 0.5669 0.5350 0.5676
CodeLlama-7b-hf 0.4979 0.5079 0.5269 0.5172 0.5442
Accuracy: As shown in Table 8, UniXcoder-base achieves the highest accuracy of 0.7172, followed by GraphCodeBERT-base at 0.6517. In contrast, CodeLlama-7b-hf yields the lowest accuracy among the fine-tuned CodeLMs at 0.4979. For reference, the inference-only GPT-4o-mini baseline achieves an accuracy of 0.5196, indicating that fine-tuned CodeLMs effectively enhance the overall prediction accuracy and significantly outperform the general LLM under zero-shot settings in the CLB detection task.
Precision: The precision of the fine-tuned models varies, with UniXcoder-base achieving the highest precision at 0.6839, followed by CodeT5-base at 0.6350 and CodeT5p-220M at 0.6148. These results suggest that fine-tuning enables the models to effectively identify positive instances while controlling false positives. In comparison, GPT-4o-mini achieves a precision of 0.5190, which is surpassed by almost all fine-tuned CodeLMs.
Recall: Recall reflects the model’s ability to capture positive instances. Several fine-tuned CodeLMs demonstrate strong recall, such as CodeT5p-770M (0.8779) and GraphCodeBERT-base (0.8241), indicating high sensitivity in recognizing CLBs. Conversely, Qwen2.5_Coder-7B and CodeLlama-7b-hf show lower recall scores of 0.5669 and 0.5269, respectively. The GPT-4o-mini baseline achieves a recall of 0.5357, further highlighting that task-specific adaptation generally leads to higher sensitivity in bug detection.
F1 score: UniXcoder-base achieves the highest F1 score of 0.7407, demonstrating an excellent balance between precision and recall. GraphCodeBERT-base also performs well with an F1 score of 0.7029. In comparison, the F1 score of GPT-4o-mini is 0.5272, which is consistently lower than those of almost all fine-tuned CodeLMs, highlighting the effectiveness of task-specific fine-tuning for CLB detection.
AUC: The AUC metric further confirms the discriminative capability of the fine-tuned CodeLMs. UniXcoder-base reaches the highest AUC of 0.8088, indicating a strong ability to distinguish between positive and negative samples, particularly in challenging scenarios such as semantically similar code snippets. Similarly, the AUC of GPT-4o-mini is 0.5196, which is notably lower than those of the fine-tuned CodeLMs, indicating limited discriminative capability under inference-only settings.